The car at rest moves as 9 km/hr in 2hrs. Find the acceleration.


1
Expert's answer
2021-05-23T16:33:52-0400
"a=\dfrac{v-v_0}{t},""a=\dfrac{9\ \dfrac{km}{h}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h}{3600\ s}-0}{7200\ s}=3.5\cdot10^{-4}\ \dfrac{m}{s^2}."